a nut store normally sells cashews for $9.00 per pound and almonds for $3.50 per pound. but at the end of the month the almonds had not sold well, so in order to sell 60 pounds of almonds the manager decided to mix the 60 pounds of of almonds with some cashews and sell the mixture for $7.50per pound. How many pounds of chashews should be mixed with the almonds to ensure no change in the profit?
Answer by mananth(16946)

a nut store normally sells cashews for $9.00 per pound and almonds for $3.50 per pound. but at the end of the month the almonds had not sold well, so in order to sell 60 pounds of almonds the manager decided to mix the 60 pounds of of almonds with some cashews and sell the mixture for $7.50per pound. How many pounds of chashews should be mixed with the almonds to ensure no change in the profit?
cashew $ 9 /lb
almonds $ 3.50 / lb
Let cashews mixed be x lbs
almonds mixed is 60 lbs
9x+3.50*60 = 7.5(x+60)
9x+210=7.5x+450
1.5x= 240
x= 240/1.5 = 160 lbs of cashews